1) Use the Income Statement and Balance Sheet to calculate the most important 2014 working capital ratios. Hampton Bay 2013 and 2014 Balance Sheets (S millions) 2013 2014 2013 2014 Current assets Current liabilities 74 165 393 632 Accounts payable Cash Accounts receivable Inventory 98 178 422 698 312 $ 231 543 $ 531 $ 344 196 540 457 Notes payable Total Total Long-term debt Owners equity Fixed assets Common stock and paid-in surplus Retained earnings 500 $ 1,799 2,299 S 550 2,041 2,591 Net plant and equipment 2,731 2,880 Total Total assets 3,363 $ 3,578 Total liabilities and equity 3,373 $ 3,588 S Hampton Bay 2014 Income Statement Sales Cost of goods sold Depreciation Earnings before interest and taxes Interest paid Taxable income Taxes Net income 2,311 1,244 276 791 141 650 187 463 Dividends Additions to retained earnings 121 342

solve

a. return on equity
b. Cash collections from customers
c. cash paid to suppliers
Answer #1

a)Return on equity = net income / Average equity

= 463 /[(2299+2591)/2]

    = 463 /2445

    = .1894 or 18.94%

b)cash collection from customers = Beginning receivables + sales -ending receivables

      =165+2311-178

        = 2298

**assuming all sales are on credit

c) cash paid to suppliers = Beginning accounts payables + purchase -ending accounts payables

              = 312+1273 - 344

               = 1241

**Purchase =cost of goods sold + ending inventory - beginning inventory

=1244+422-393

   = 1273